OpenStudy (anonymous):

5x + 10y = -20, solve for y

OpenStudy (anonymous):

My answer isa. y = -½ x - 2

OpenStudy (anonymous):

y-intercept = -4/2 = -2

OpenStudy (anonymous):

y=-2

OpenStudy (alexandervonhumboldt2):

first subtract 5x from both sides: 5x-5x+10y=-20-5x 10y=-20-5x divide both sides by 10 10y/10=(-20-5x)/10 y=-2-1/2x y=-1/2x-2

OpenStudy (confxsedaf):

you're correct @ninamae05

OpenStudy (alexandervonhumboldt2):

you are correct @ninamae05